# Multi-Prompting Decoder Helps Better Language Understanding

### v1

@@ Abstract
Recent Pre-trained Language Models (PLMs) usually only provide users with the inference APIs, namely the emerging Model-as-a-Service (MaaS) setting. To adapt MaaS PLMs to downstream tasks without accessing their parameters and gradients, some existing methods focus on the output-side adaptation of PLMs, viewing the PLM as an encoder and then optimizing a task-specific decoder for decoding the output hidden states and class scores of the PLM. Despite the effectiveness of these methods, they only use a single prompt to query PLMs for decoding, leading to a heavy reliance on the quality of the adopted prompt. In this paper, we propose a simple yet effective Multi-Prompting Decoder (MPD) framework for MaaS adaptation. The core idea is to query PLMs with multiple different prompts for each sample, thereby obtaining multiple output hidden states and class scores for subsequent decoding. Such multi-prompting decoding paradigm can simultaneously mitigate reliance on the quality of a single prompt, alleviate the issue of data scarcity under the few-shot setting, and provide richer knowledge extracted from PLMs. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated decoding for class scores. Extensive experiments demonstrate that our method achieves new state-of-the-art results on multiple natural language understanding datasets under the few-shot setting. All the code and data of this work can be obtained at https://github.com/czlll/MPD.
Index Terms:
Model-as-a-Service, few-shot learning.
I Introduction
@@
Pre-trained Language Models (PLMs) have recently achieved remarkable performance on various downstream language understanding tasks with the “pre-training and fine-tuning” paradigm [1]. However, as the scale of PLMs continues to grow, fine-tuning the full model for each downstream task has become computationally expensive and deployment-inefficient. In light of this, Model-as-a-Service (MaaS) [2, 3] has emerged as a more practical model deployment scheme, allowing pre-trained models to serve downstream tasks by providing inference APIs. Within the MaaS scheme, users can access the outputs of PLMs (e.g., hidden states, class scores, or textual output) through APIs [^1] Currently, some famous MaaS services include GPT-3.5 Turbo, which offers APIs for accessing class scores and text output, as well as text-embedding-3-small and text-embedding-3-large, which provide APIs for accessing embedding representations (i.e., hidden states) of input text., but cannot access the parameters and gradients of PLMs, presenting a challenge in effectively adapting PLMs to downstream tasks. As shown in Table I, MaaS belongs to the category of grey-box PLM adaptation methods, which, unlike white-box adaptation methods, don’t require access to model parameters.
TABLE I: Comparison of existing PLM adaptation methods. Compared to kNN-Adapter, MaaS focuses on downstream task adaptation and does not require retrieval corpus.
Methods w/o Model Parameters w/o Hidden State w/o Token Probabilities w/o Retrieval Corpus
   White-Box LLM Fine-Tuning: full access to PLM
Fine-Tuning
Adapter [4]
Prefix-Tuning [5]
LoRA [6]
   Grey-Box LLM Adaptation: can not access the parameters of PLM
kNN-Adapter [7]
MaaS [3]
@@
To adapt MaaS PLMs to downstream tasks, existing methods have primarily explored both the input-side and the output-side adaptation. On the input-side adaptation, researchers focus on utilizing gradient-free methods to heuristically search a good continuous [3, 8] or discrete prompt [2, 9] for the target task, such as using evolutionary algorithm and reinforcement learning. However, input-side methods need to query PLM thousands of times for optimization, resulting in significant time overhead and the optimization process is very difficult due to the huge search space and the lack of gradients. On the output-side adaptation, researchers [10, 11] treat the PLM as an encoder and locally train a task-specific decoder to post-process PLM’s outputs. Compared to the input-side adaptation, the output-side adaptation methods only query PLM a few times and can benefit from gradient for optimization, often resulting in better performance.
Despite the effectiveness of these output-side adaptation methods, they usually only adopt a single prompt to query PLMs for a certain downstream task, thus often struggling with the performance variations brought by using different prompts [12, 13]. As shown in Figure 1, on all four datasets, the maximum performance fluctuations in accuracy can exceed 8% when using different prompts. Therefore, the quality of prompt is crucial, and relying on only one prompt easily exposes them to performance risks of encountering a low-quality query.
@@
Based on the above analysis, in this paper, we propose a novel output-side PLMs adaptation framework for few-shot classification, named Multi-Prompting Decoder (MPD). This MPD framework can decode the outputs of PLMs, including hidden states and class scores, to yield classification results. In contrast to previous methods, MPD adopts a novel multi-prompting decoding paradigm, which queries PLMs with multiple prompts for a sample, thereby obtaining more representation information for decoding. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated multi-prompting decoding for class scores. Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks.
In summary, the major contributions of this paper are:
  • We propose an output-side PLMs adaptation framework, multi-prompting decoder, for few-shot classification.
  • We design two multi-prompting decoding strategies: multi-prompting decoding with optimal transport and calibrated multi-prompting decoding to decode hidden states and class scores respectively.
  • Extensive experiments demonstrate that our method achieves new state-of-the-art results under the few-shot setting on multiple natural language understanding tasks, including sentiment analysis, topic classification, and natural language inference. Furthermore, our framework is demonstrated to be lightweight, efficient, and applicable to PLMs of different architectures (e.g., encoder only, decoder only, and encoder-decoder).
The rest of the article is organized as follows. Section II provides a comprehensive review of related work on prompt tuning, MaaS adaptation, parameter-efficient tuning, and optimal transport. Section III describes the preliminaries about MaaS adaptation and optimal transport. Section IV describes the technical details of our proposed MPD. Experimental settings, comparison, and analyses are presented in Section V. Finally, we conclude our work in Section VI.
II Related Work
Our work focuses on how to efficiently adapt PLMs to downstream tasks. We review four lines of research: prompt tuning (data efficiency), MaaS adaptation methods, parameter-efficient tuning (parameter efficiency), and optimal transport respectively.
II-A Prompt Tuning
@@
Prompt tuning aims to reduce the gap between PLM pre-training and fine-tuning via predicting [MASK] token in the prompt for data-efficiency and achieves great performance in zero-shot and few-shot learning [1]. Prompt-based tuning is originally applied in knowledge probing [14, 15], and it has been adopted broadly in NLP [16, 17, 18, 19]. Since prompt has a large impact on the performance of prompt tuning, many studies have focused on learning a good prompt to alleviate the prompt engineering, including discrete prompt [20, 21, 22] and continuous prompt [23, 24, 5, 25, 26]. However, optimizing prompts is difficult [25] and sometimes computationally expensive and not interpretable. Our work focuses on output-side adaptation and uses manual prompts to make PLM adaptable to downstream tasks.
II-B MaaS Adaptation
@@
In MaaS settings, users can access the outputs of PLMs (e.g., hidden states, class scores, or textual output) via APIs, but cannot access the parameters and gradients of PLMs. Thus, some works try to adapt these models to downstream tasks without accessing the model parameters and the gradients [3, 2].
@@
Some input-side adaptation work about MaaS uses reinforcement learning [2, 9] or derivative-free optimization [3, 8] to find the optimal prompt. For the first type of method, Diao et al. [2] first use a variance-reduced policy gradient algorithm to estimate the gradient of the prompt token distribution. Deng et al. [9] use a policy network and reward function to generate and optimize discrete prompts. TEMPERA [27] constructs query-dependent prompts through test-time prompt editing and formulates this as a reinforcement learning problem. DP2[28] first designs a prompt generation strategy through multi-round dialogue alignment on GPT-4 and prompt evaluation metric SUE. Then, DP2O constructs a reinforcement learning framework based on policy gradients to match the prompts to inputs optimally. For the second type of method, BBT [3] adopts a covariance matrix adaptation evolution algorithm to optimize continuous prompt. BBTv2 [8] further prepends continuous prompts to every layer of PLM and proposes a divide-and-conquer gradient-free algorithm to optimize them. GDFO [29] uses a prompt generator to generate an extra vector as input based on the BBT framework. However, searching in a large space of prompts is inefficient and difficult to optimize.
@@
Inspired by this, output-side adaptation methods use manual prompts and train task-specific decoder to post-process PLM’s outputs. PromptBoosting [10] constructs a series of weak learners, learns the corresponding verbalizer, and uses AdaBoost algorithm to ensemble them. DecT [11] extracts text feature and optimizes class prototypes as hyperspheres with a radius parameter.
II-C Parameter-Efficient Tuning
Parameter-Efficient Tuning (PET) aims to adjust only a small number of parameters while keeping most pre-trained parameters frozen. Houlsby et al. [4] first add two lightweight adapter modules to every layer. Afterwards, prefix tuning [5] prepends multiple trainable prefix vectors to the keys and values of the multi-head attention at every layer and LoRA [6] injects trainable low-rank matrices into every layer to approximate the weight updates. Besides, Lester et al. [25] prepend continuous prompt tokens to the input layer, Zaken et al. [30] only optimize the bias vectors, and He et al. [31] present a unified framework that establishes connections between those PET methods. Despite their success with few parameter updates, they still require model gradients to update the small portion of parameters, which are unavailable in the MaaS setting.
II-D Optimal Transport
@@
OT theory [32] was originally used to study the movement of many items from one place to another with minimal cost and provides a distance measure method between two distributions in a closed way. Due to the excellent properties of the distance measure, OT is used in many areas, such as topic model [33], legal case matching [34], vision-language models [35], and image classification [36]. In this paper, we use OT to align text features obtained using multiple prompts and multiple class prototypes.
III Preliminaries
We present two preliminaries that are relevant to our work, i.e., model-as-a-service adaptation and optimal transport.
III-A MaaS Adaptation
@@
MaaS adaptation aims to adapt the PLMs to downstream tasks based on a given few-shot training set. Specifically, the setting needs to correctly predict the label $y\in\{1,\dots,N\}$ for test sample $x$ based on a given few-shot training set $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{NK}$ and PLM $\mathcal{M}$, where $N$ is the number of classes and each class has $K$ training samples. In the MaaS setting, we can only query the PLM model $\mathcal{M}$ with input $x$ and get corresponding outputs (e.g., hidden states and class scores) without accessing the model parameters and the gradients.
@@
III-B Optimal Transport

### v2

@@ Abstract
Recent large Pre-trained Language Models (PLMs) usually only provide users with the inference APIs, namely the emerging Model-as-a-Service (MaaS) setting. To adapt MaaS PLMs to downstream tasks without accessing their parameters and gradients, some existing methods focus on the output-side adaptation of PLMs, viewing the PLM as an encoder and then optimizing a task-specific decoder for decoding the output hidden states and class scores of the PLM. Despite the effectiveness of these methods, they only use a single prompt to query PLMs for decoding, leading to a heavy reliance on the quality of the adopted prompt. In this paper, we propose a simple yet effective Multi-Prompting Decoder (MPD) framework for MaaS adaptation. The core idea is to query PLMs with multiple different prompts for each sample, thereby obtaining multiple output hidden states and class scores from PLMs for subsequent decoding. Such multi-prompting decoding paradigm can simultaneously mitigate reliance on the quality of a single prompt, alleviate the issue of data scarcity under the few-shot setting, and provide richer knowledge extracted from PLMs. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated decoding for class scores. Extensive experiments demonstrate that our method is effective on multiple natural language understanding datasets under the few-shot setting.
Multi-Prompting Decoder Helps Better Language Understanding
Zifeng Chengthanks:   Both authors contributed equally to this research., Zhaoling Chen11footnotemark: 1, Zhiwei Jiangthanks:   Corresponding author., Yafeng Yin, Cong Wang, Shiping Ge, Qing Gu National Key Laboratory for Novel Software Technology, Nanjing University {chengzf,zhaolingchen}@smail.nju.edu.cn  {jzw,yafeng}@nju.edu.cn {cw,shipingge}@smail.nju.edu.cn   guq@nju.edu.cn
1 Introduction
@@
Pre-trained Language Models (PLMs) have recently achieved remarkable performance on various downstream language understanding tasks with the “pre-training and fine-tuning” paradigm Liu et al. (2023). However, as the scale of PLMs continues to grow, fine-tuning the full model for each downstream task has become computationally expensive and deployment-inefficient. In light of this, Model-as-a-Service (MaaS) Diao et al. (2022); Sun et al. (2022b) has emerged as a more practical model deployment scheme, allowing pre-trained models to serve downstream tasks by providing inference APIs. Within the MaaS scheme, users can access the outputs of PLMs (e.g., hidden states, class scores, or textual output) through APIs [^1] Currently, some famous MaaS services include GPT-3.5 Turbo, which offers APIs for accessing class scores and text output, as well as text-embedding-3-small and text-embedding-3-large, which provide APIs for accessing embedding representations (i.e., hidden states) of input text., but cannot access the parameters and gradients of PLMs, presenting a challenge in effectively adapting PLMs to downstream tasks.
To adapt MaaS PLMs to downstream tasks, existing methods have primarily explored both the input-side and the output-side adaptation. On the input-side adaptation, researchers focus on utilizing gradient-free methods to heuristically search a good continuous Sun et al. (2022b, a) or discrete prompt Diao et al. (2022); Deng et al. (2022) for the target task, such as using evolutionary algorithm and reinforcement learning. However, input-side methods need to query PLM thousands of times for optimization, resulting in significant time overhead and the optimization process is very difficult due to the huge search space and the lack of gradients. On the output-side adaptation, researchers Hou et al. (2023); Cui et al. (2023) treat the PLM as an encoder and locally train a task-specific decoder to post-process PLM’s outputs. Compared to the input-side adaptation, the output-side adaptation methods only query PLM a few times and can benefit from gradient for optimization, often resulting in better performance.
@@
Despite the effectiveness of these output-side adaptation methods, they usually only adopt a single prompt to query PLMs for a certain downstream task, thus often struggling with the performance variations brought by using different prompts Lu et al. (2022); Chen et al. (2023b). As shown in Figure 1, on all four datasets, the maximum performance fluctuations in accuracy can exceed 8% when using different prompts. Therefore, the quality of prompt is crucial, and relying on only one prompt easily exposes them to performance risks of encountering a low-quality query.
@@
Based on the above analysis, in this paper, we propose a novel output-side PLMs adaptation framework for few-shot classification, named Multi-Prompting Decoder (MPD). This MPD framework can decode the outputs of PLMs, including hidden states and class scores, to yield classification results. In contrast to previous methods, MPD adopts a novel multi-prompting decoding paradigm, which queries PLMs with multiple prompts for a sample, thereby obtaining more representation information for decoding. Specifically, we propose two decoding strategies: multi-prompting decoding with optimal transport for hidden states and calibrated multi-prompting decoding for class scores. Extensive experiments demonstrate that our method is effective under the few-shot setting on multiple natural language understanding tasks, including sentiment analysis, topic classification, and natural language inference.
2 Related Work
@@
Prompt Tuning Prompt tuning aims to reduce the gap between PLM pre-training and fine-tuning via predicting [MASK] token in the prompt for data-efficiency and achieves great performance in zero-shot and few-shot learning Liu et al. (2023). Since prompt has a large impact on the performance of prompt tuning, many studies have focused on finding a good prompt, including discrete prompt Jiang et al. (2020); Shin et al. (2020); Gao et al. (2021), continuous prompt Liu et al. (2021b); Qin and Eisner (2021); Li and Liang (2021); Lester et al. (2021); Zhang et al. (2022), prompt selection Sorensen et al. (2022); Chen et al. (2023b), and so on.
@@
MaaS Adaptation Some PLMs, such as GPT-3 Brown et al. (2020), are released as a service in the cloud. Some works try to adapt these models to downstream tasks without accessing the model parameters and the gradients Sun et al. (2022b); Diao et al. (2022).
@@
Some work about MaaS uses reinforcement learning (RL) Diao et al. (2022); Deng et al. (2022) or derivative-free optimization Sun et al. (2022b, a) to find the optimal prompt. For the first type of method, Diao et al. (2022) first use a variance-reduced policy gradient algorithm to estimate the gradient of the prompt token distribution. Deng et al. (2022) use a policy network and reward function to generate and optimize discrete prompts. For the second type of method, BBT Sun et al. (2022b) adopts a covariance matrix adaptation evolution algorithm to optimize continuous prompt. BBTv2 Sun et al. (2022a) further prepends continuous prompts to every layer of PLM and proposes a divide-and-conquer gradient-free algorithm to optimize them. GDFO Han et al. (2023) uses a prompt generator to generate an extra vector as input based on the BBT framework. TEMPERA Zhang et al. (2023) constructs query-dependent prompts through test-time prompt editing and formulates this as an RL problem. DP2Li et al. (2023) first designs a prompt generation strategy through multi-round dialogue alignment on GPT-4 and prompt evaluation metric SUE. Then, DP2O constructs RL framework based on policy gradients to match the prompts to inputs optimally.
@@
However, searching in a large space of prompts is inefficient and difficult to optimize. Inspired by this, PromptBoosting Hou et al. (2023) constructs a series of weak learners, learns the corresponding verbalizer, and uses AdaBoost algorithm to ensemble them. DecT Cui et al. (2023) is an output-side adaptation framework and optimizes class prototypes as hyperspheres with a radius parameter.
It’s worth noting that parameter-efficient tuning Houlsby et al. (2019); Hu et al. (2022); Zaken et al. (2022); He et al. (2022) requires gradients of PLMs to update the small portion of parameters, which are unavailable in the MaaS setting.
Figure 2: Illustration of the multi-prompting decoding with optimal transport.
@@
Optimal Transport OT theory Monge (1781) was originally used to study the movement of many items from one place to another with minimal cost and provides a distance measure method between two distributions in a closed way. Due to the excellent properties of the distance measure, OT is used in many areas, such as topic model Zhao et al. (2021a), vision-language models Chen et al. (2023a), and image classification Liu et al. (2021a).
3 Preliminaries
@@
MaaS adaptation aims to correctly predict the label $y\in\{1,\dots,N\}$ for test sample $x$ based on a given few-shot training set $\mathcal{D}_{\text{train}}=\{(x_{i},y_{i})\}_{i=1}^{NK}$ and PLM $\mathcal{M}$, where $N$ is the number of classes and each class has $K$ training samples. In the MaaS setting, PLM $\mathcal{M}$ is a black-box inference API without accessing the model parameters and the gradients. Therefore, we can only query the model with input $x$ and get corresponding outputs, including hidden states and class scores.
@@
4 Methodology
We introduce a novel output-side MaaS adaptation framework, Multi-Prompting Decoder (MPD), for few-shot classification. This framework can decode the outputs of PLMs, including hidden states and class scores, to yield classification results. In contrast to previous decoding frameworks that query PLMs only once for a sample, we propose multi-prompting decoding paradigm, which queries PLMs with multiple prompts for a sample, thereby obtaining more representation information for decoding. Specifically, for decoding the PLMs’ output hidden states, we design a multi-prompting decoding strategy based on optimal transport. For decoding the PLMs’ output class scores, we additionally design a calibrated multi-prompting decoding strategy. These two decoding strategies are jointly used in the final decoding process.
4.1 Multi-Prompting Decoding with Optimal Transport
In our multi-prompting paradigm, each sample has multiple types of hidden states corresponding to multiple prompts. Building a unified classifier for these hidden states neglects the characteristics of each prompt, leading to a crude solution. Building multiple classifiers specific to each type of hidden state is also a sub-optimal solution, which will lead to an ensemble of multiple classifiers, and these classifiers cannot benefit from each other during training.
To fully exploit the advantages of the multi-prompting decoding paradigm, we design a specific multi-prompting decoder for hidden states. The core idea is to establish multiple prototypes for each class to capture the class characteristics specific to different prompts. Subsequently, through optimal transport, we identify the best matching between the sample representations and the prototypes, thereby achieving classification. Considering the limited number of training samples available in the few-shot setting, the decoder is lightweight, comprising only a linear layer and several learnable class prototypes.
4.1.1 Querying with Multiple Prompts
Given a text $x_{i}$, we first use a series of templates $\mathcal{T}_{1},\cdots,\mathcal{T}_{P}$ with a [MASK] token to enclose the text and then query $\mathcal{M}$ with template-wrapped inputs $\mathcal{T}_{1}(x_{i})$, $\cdots$, $\mathcal{T}_{P}(x_{i})$ to get a series of initial text representations $\mathbf{h}_{i,1},\cdots,\mathbf{h}_{i,P}$. Specifically, we take hidden states of the final layer at the [MASK] position as the initial text representations extracted by PLM $\mathcal{M}$. Then, we use a linear layer parameterized by $\mathbf{W}$ to project the initial representations to get text representations for classification:
$\mathbf{v}=\mathbf{Wh}$ (1)
4.1.2 Features-Prototypes Matching via OT
Since our framework contains multiple text representations and class prototypes, OT is a proper solution to better map multiple text representations and multiple prototypes of each class. OT aims to find the optimal transport plan $\bm{T}\in\mathbb{R}^{P\times Q}$ that represents the fine-grained matching flow between text representations and class prototypes, where $P$ denotes the number of prompts and $Q$ denotes the number of prototypes. The optimal transport plan assigns a higher weight to the pair of text representation and prototype that are close to each other. We use Sinkhorn’s algorithm Cuturi (2013) to optimize the transport plan. Specifically, we fix text representations and class prototypes to optimize the transport plan. The details of OT and Sinkhorn’s algorithm are shown in Appendix A.
4.1.3 Classifying with OT Scores
After getting the optimal transport plan, the optimal transport score between text $x_{i}$ and class $k$ can be computed according to the optimal transport plan and similarity matrix. Specifically,
$\text{Dec}(x_{i},k)=\sum_{m=1}^{P}\sum_{n=1}^{Q}\bm{T}^{i,k}_{m,n}\text{sim}(\mathbf{V}_{i,m},\mathbf{R}_{k,n})$ (2)
where Dec(,) denotes the optimal transport score between text $x_{i}$ and class $k$, $\bm{T}^{i,k}_{m,n}$ denotes the ($m$,$n$)-th element in the optimal transport plan between text $x_{i}$ and class $k$, sim(,) denotes cosine similarity, $\mathbf{V}_{i,m}$ is the text representations of $\mathcal{T}_{m}(x_{i})$, and $\mathbf{R}_{k,n}$ is the $n$-th prototype representation of the $k$-th class.
Finally, we use the softmax function over optimal transport scores between text representations and all classes to get the predicted probability distribution, and cross-entropy loss to optimize the parameters. Specifically,
The whole training flow is shown in Algorithm 1.
4.2 Calibrated Multi-Prompting Decoding
We also decode the PLMs’ output class scores as supplementary for the hidden states decoding to provide additional prior knowledge of the PLMs. Specifically, we first expand the set of label words for the verbalizer, then calibrate the class scores for each prompt separately, and average the calibrated class scores of all prompts. Finally, we combine OT scores and calibrated class scores for joint decoding.
4.2.1 Label Words Expansion
We first use a simple and effective way to expand the label words. Given the prediction layer of MLM $\mathbf{W}_{mlm}\in\mathbb{R}^{|\mathcal{V}|*d}$, we can treat each item as a representation of a word. Then, given a label word, we expand a new set of label words S based on cosine similarity between words and use the softmax function over cosine similarity to assign a weight to each element in the set. It is worth noting that due to the large number of open-sourced PLMs, getting an expanded set of label words is easy.
Figure 3: Illustration of the calibrated multi-prompting decoding.
4.2.2 Calibrating Class Scores
Predictions of PLM are biased because PLM tends to predict tokens that are common in its pre-training distribution Zhao et al. (2021b). Thus, we first calibrate PLM on a given prompt and then use calibrated class scores for classification.
Specifically, we use a template to wrap to empty input $x_{e}=$“” and query the model with $\mathcal{T}_{j}(x_{e})$ to obtain the predictions of label words $\mathbf{s}_{ej}\in\mathbb{R}^{N|\emph{S}|}$ to calibrate. Then, given a prediction $\mathbf{s}_{ij}$ for sample $x_{i}$ wrapped by prompt $\mathcal{T}_{j}$, we can calibrate $\mathbf{s}_{ij}$ by
where $\mathbf{\bar{s}}_{ij}\in\mathbb{R}^{N|\emph{S}|}$ is calibrated prediction over the expanded set. We can see that the calibrated predictions for context-independent inputs such as empty input become uniformly distributed.
Then, we use the weight vector for each class to aggregate the predictions of each element in the corresponding set to form the calibrated class scores $\mathbf{\hat{s}}_{ij}\in\mathbb{R}^{N}$. Finally, we average the class scores of multiple prompts to get averaged calibrated class scores $\mathbf{\hat{s}}_{i}$ for classification.
$\mathbf{\hat{s}}_{i}=\frac{\sum_{j=1}^{P}\mathbf{\hat{s}}_{ij}}{P}$ (5)
4.2.3 Joint Decoding
$\hat{y}_{i}=\hat{y}^{OT}_{i}+\beta\mathbf{\hat{s}}_{i}$ (6)
$N$ Method SST2 IMDB Yelp AG DB Yahoo RTE SNLI MNLI-m/mm Avg.
Table 1: Experiment results for MaaS adaptation methods. The results (i.e., average accuracy and standard deviation (%)) over 5 runs. The best results are in bold.
5 Experiments
5.1 Datasets
We conduct experiments on several common natural language understanding tasks including sentiment analysis, topic classification, and natural language inference (NLI). For sentiment analysis, we choose SST2 Socher et al. (2013), IMDB Maas et al. (2011), and Yelp Zhang et al. (2015). For topic classification, we choose AG’s News Zhang et al. (2015), DBPedia Zhang et al. (2015), and Yahoo Zhang et al. (2015). For NLI, we choose RTE Dagan et al. (2005), SNLI Bowman et al. (2015), and MNLI Williams et al. (2018). The statistics of datasets are shown in Table 5.
We also randomly sample $N$ = 1, 4, 16 samples for each class from the original training set to construct the few-shot training set. Following Cui et al. (2023) and Hou et al. (2023), we use the original validation set for evaluation on the GLUE (i.e., SST, RTE, and MNLI) and SNLI datasets.
5.2 Baselines
We compare with some strong MaaS methods. In-context learning (ICL) Brown et al. (2020) concatenates some text-label pairs before the test samples. BBT Sun et al. (2022b) optimizes soft prompt tokens with a derivative-free evolutionary algorithm. BBTv2 Sun et al. (2022a) further prepends continuous prompts to every layer and uses a divide-and-conquer algorithm to optimize them. RLPrompt Deng et al. (2022) optimizes discrete prompts with RL and adopts soft Q-learning to find the optimal prompt. PromptBoosting Hou et al. (2023) sequentially learns multiple weak learners and adopts AdaBoost to ensemble them. DecT Cui et al. (2023) extracts text feature and optimizes class prototypes as hyperspheres with an additional radius parameter.
5.3 Implementation Details
[…]
Method Trainable Param (K) Query Number SST2 AG
Acc Training Time (s) Acc Training Time (s)
ICL 0 0 71.5 0 64.4 0
BBT 0.5 8,000 89.6 1619 87.8 5228
BBTv2 12 8,000 90.3 1624 85.3 6546
RLPrompt 3100 12,000 87.0 82286 80.2 99947
PromptBoosting 0.4 10 87.6 190 85.2 398
DecT 130 1 91.0 1.4 86.4 2.3
Ours 132 3 91.9 3.5 87.9 6.7
Table 2: Efficiency comparison for MaaS methods under 16-shot setting on two datasets.
5.4 Results
Table 1 shows the performance on 9 datasets under the few-shot setting. We can see that our model achieves state-of-the-art performance. Specifically, compared to DecT, our approach achieves 1.5%, 3.4%, and 2.3% average performance improvement under 1,4,16-shot settings. Among all settings, our model performs the best in 27 out of the 30. This shows our model is effective and the improvement is consistent. It is worth noting that our model performs well on both simple (i.e., sentiment analysis and topic classification) and difficult tasks that require semantic understanding (i.e., NLI). This shows that the improvement is task-independent.
By observing other baseline methods, ICL does not achieve performance gains, as the shot number increases. This is due to the length limitation of the PLM. BBTv2 exceeds BBT and RLPrompt. This shows that inserting continuous prompts to every layer of the PLM is effective. PromptBoosting and DecT usually achieve better performance than BBT and BBTv2. This is because these two methods do not need to search prompts in a large space and can benefit from gradients.
$N$ Method SST2 IMDB Yelp AG DB Yahoo RTE SNLI MNLI-m/mm Avg.
Table 3: Experiment results for more training data. The results over 5 runs. Fine-tuning method requires gradients of the model to update the model.
Model Setting Average Accuracy
1 4 16
Ours 71.5 75.2 77.8
w/o Class Scores 57.8 69.6 76.5
w/o OT Scores 68.0 68.0 68.0
w/o Expansion 70.7 74.8 77.7
Table 4: The ablation study of our model on all datasets. We run each experiment over 5 random seeds.
5.5 Efficiency Comparison
[…]
5.6 Performance on More Training Data
We conduct experiments with a larger training set (i.e., $N=$ 64 and 256) to explore the scalability of our model beyond the few-shot setting. We compare our model with DecT and a strong baseline fine-tuning. Fine-tuning requires gradients of PLM to update full parameters of PLM.
As shown in Table 3, firstly, our model and DecT both achieve performance improvement when the shot number increases. This shows that these two methods can still fit larger datasets even with a small number of tunable parameters. In addition, as the shot number increases, the average standard deviation of two models decreases. Secondly, our model also exceeds DecT when $N$ is large. This shows that the improvement is consistent. It is worth noting that the gap between the two methods is narrowing as the shot number increases. Thirdly, our method exceeds fine-tuning in the 64-shot setting and is exceeded by fine-tuning in the 256-shot setting. This is because our method cannot update PLM parameters, it can only update a small portion of parameters. Thus, when the shot number is larger, our model may not be able to learn more knowledge. Finally, similar to DecT, our model does not show significant performance differences compared to fine-tuning in simple tasks (i.e., sentiment analysis and topic classification). However, there are significant differences in harder tasks (i.e., SNLI). This suggests that harder tasks require more parameters to learn corresponding semantics.
6 Model Analysis
6.1 Ablation Study
We conduct ablation study to validate the effectiveness of each component. As shown in Table 4, firstly, all components achieve performance improvement. Secondly, the performance of w/o Class Scores (i.e., only using OT scores) improves rapidly as the shot number increases. This shows that shot number has a significant effect on OT scores. Thirdly, w/o OT Scores (i.e., only averaging calibrated class scores of multiple prompts) also achieves good performance, especially in the 1-shot setting. This shows that averaging the class scores of multiple prompts is also a strong baseline under 1-shot setting. Finally, label words expansion achieves more significant performance gains when the shot number is small. This is because the model relies more on class scores when the shot number is small. The results of label words expansion are shown in Appendix E.
6.2 Effects of Prompts
We explore the effects of prompts by changing the combination of prompts in Figure 4. Specifically, we use different prompts and change the number of prompts. The five templates are shown in Table 13.
Firstly, as the prompt number increases, performance will be better and more stable. When the number of prompts exceeds 3, the improvement is marginal. Secondly, there will be less performance fluctuation with multiple prompts compared to using a single prompt Specifically, when we use one/two/three/four prompts, the change in performance is 1.7%/0.3%/0.2%/0.2%. This shows that our method is robust across multiple prompts. Thirdly, the good performance of a prompt’s class scores does not mean that using it can achieve better performance. This shows that there is a gap between the representation of hidden states and class scores.
Figure 5: Effects of $\beta$ over 5 runs on two datasets.
6.3 Effects of $\beta$
Hyper-parameter $\beta$ controls the ratio of OT scores and class scores. We explore the effects of $\beta$ on two datasets in Figure 5. Firstly, our model needs a bigger $\beta$ under 1-shot setting. This is because the amount of data is small and the model relies heavily on class scores for prediction rather than OT scores. Secondly, as the shot number increases, the effect of $\beta$ on the results diminishes. This is because as the shot number increases, the model’s capability is enhanced and relies more on the OT scores for prediction. Thirdly, as the shot number increases, the standard deviation becomes smaller. This shows that the OT scores are more stable when the training size increases.
Figure 6: Effects of $Q$ over 5 runs on two datasets.
6.4 Effects of the Number of Prototypes
We explore the effects of the number of prototypes on two datasets in Figure 6. We can see that using multiple prototypes tends to outperform using one. This shows that using multiple prototypes is valid in our multiple prompts setting. As the shot number increases, a large $Q$ usually performs better. This is because when the shot number increases, the model needs a large $Q$ to capture subtle differences between multiple texts of the same class.
7 Conclusion
In this paper, we propose the multi-prompting decoder framework for output-side MaaS adaptation on few-shot tasks. To achieve better decoding performance, we design two decoding strategies to decode PLMs’ output hidden states and class scores, respectively. Extensive experiments show that our approach is effective and efficient.
Limitations
Although our approach achieves state-of-the-art performance, it has some limitations.
The first limitation is that our approach relies on two kinds of PLMs’ outputs, namely hidden states and class scores, but these outputs are not always available for some PLMs, such as ChatGPT. More rigorous decoding, such as decoding only based on the accessing of the output tokens, is worth further exploration.
The second limitation is that we have only conducted experiments on the English datasets. The performance of our method on datasets in other languages deserves further exploration.
The third limitation is that our method relies on some manual engineering, such as the design of prompts and label words. In this paper, this problem is alleviated by using templates and label words from previous works.
The fourth limitation is that MPD may not be suitable for training with more training data. As shown in Table 1 and Table 3, the performance gap of between MPD and DecT is narrowing as the shot number $N$ increases from 4 to 256. When $N$ is 4, 16, 64, and 256, the average performance gap is 3.4, 2,3, 1.4, and 0.8, respectively. Considering that as the shot number $N$ increases, the performance improvement becomes marginal, and using multiple prompts for querying incurs additional costs, we think that our MPD is not suitable in this situation.
Acknowledgments
We would like to thank the anonymous reviewers for their insightful comments. This work is supported by the National Natural Science Foundation of China under Grants Nos. 62441225, 61972192, 62172208, 61906085. This work is partially supported by Collaborative Innovation Center of Novel Software Technology and Industrialization. This work is supported by the Fundamental Research Funds for the Central Universities under Grant No. 14380001.
Appendix A Optimal Transport